# Generalized Face Liveness Detection via De-spoofing Face Generator

### v1

@@ Abstract
Previous Face Anti-spoofing (FAS) works face the challenge of generalizing in unseen domains. One of the major problems is that most existing FAS datasets are relatively small and lack data diversity. However, we find that there are numerous real faces that can be easily achieved under various conditions, which are neglected by previous FAS works. In this paper, we conduct an Anomalous cue Guided FAS (AG-FAS) method, which leverages real faces for improving model generalization via a De-spoofing Face Generator (DFG). Specifically, the DFG trained only on the real faces gains the knowledge of what a real face should be like and can generate a “real” version of the face corresponding to any given input face. The difference between the generated “real” face and the input face can provide an anomalous cue for the downstream FAS task. We then propose an Anomalous cue Guided FAS feature extraction Network (AG-Net) to further improve the FAS feature generalization via a cross-attention transformer. Extensive experiments on a total of nine public datasets show our method achieves state-of-the-art results under cross-domain evaluations with unseen scenarios and unknown presentation attacks.
keywords:
@@ 1 Introduction
In today’s world, face recognition systems are extensively utilized for identity authentication in various scenarios. However, these systems are susceptible to presentation attacks, such as using a video or a photograph of a person to deceive the system, thereby posing a significant security threat. In order to address the security threat, researchers have constructed a diverse set of Face Anti-spoofing (FAS) datasets [1, 2, 3, 4, 5, 6, 7, 8, 9] and trained models on these datasets for face presentation attack detection. Recent works have achieved excellent performance on each of these datasets using intra-dataset evaluation protocols [10, 11, 12, 13, 14, 15, 16]. However, most existing FAS datasets are relatively small and lack data diversity. No individual dataset can cover all the different presentation attacks and backgrounds, which leads to severe domain bias for the FAS training process. As a result, models trained on a single FAS dataset often exhibit poor performance on other unseen FAS datasets [8, 17].
@@
To address this challenge, researchers adopt domain generalization (DG) techniques to enhance the model’s robustness to unseen domains. Generally, the model is trained using multiple FAS datasets simultaneously to extract a discriminative domain-invariant feature for liveness detection. Some works propose to align the extracted FAS features from different domains to make them domain indistinguishable [17, 18, 19, 20]. To further align the FAS feature, the disentanglement method is introduced to separate the domain-invariant feature from the domain-related feature  [21, 22, 23]. Meta-learning is also utilized to help the model obtain the domain-invariant feature [24, 25, 26, 27]. Although using domain generalization techniques relieves the lack of data diversity by training on multiple datasets, the capabilities of these methods are still constrained by the presentation attack types in the training FAS datasets.
@@
In this work, we find there are numerous real faces that can be easily achieved under various conditions and point out that effectively using these real faces is a key to improving the model generalization of the FAS model. Therefore, we propose an Anomalous cue Guided Face Anti-spoofing (AG-FAS) method that leverages real faces for cross-domain FAS tasks. Specifically, we train a De-spoofing Face Generator (DFG) only on the real faces to gain the knowledge of what a real face should be like. The trained DFG can then be used to generate a “real” version of the face corresponding to any given input face, achieving the “de-spoofing” of the input. Consequently, we calculate the residual between the generated image and the input, which shows the anomalous region of the input and is taken as the anomalous cue. As the DFG is trained on real faces with rich data diversity, the anomalous cue can provide robust hints about the anomaly of the input in different scenarios. In the downstream FAS task, we calculate the anomalous cue of all the images in the FAS dataset via the well-trained DFG and utilize an Anomalous cue Guided FAS feature extraction Network (AG-Net) to further obtain a more generalized FAS feature. The anomalous cue is introduced in different layers of the AG-Net via cross-attention modules to help obtain a more generalized FAS feature. Moreover, as a Plug-and-Play feature extraction module, the proposed AG-Net can be combined with most previous DG-based FAS methods to further improve the model performance across different domains. The experimental results show that the incorporation of this anomalous cue effectively enhances the performance of the model in unseen scenarios and unknown presentation attacks.
@@
  1. [1.]
  2. 1.
    We propose an Anomalous cue Guided Face Anti-spoofing (AG-FAS) method, which leverages additional real faces for improving model generalization on cross-domain FAS tasks.
  3. 2.
    We obtain an anomalous cue through a De-spoofing Face Generator (DFG) trained on the real faces. The DFG can generate a “real” face version corresponding to the input when taking the input face identity as guidance. The residual of the generated “real” image and the input is then taken as the anomalous cue for face liveness detection.
  4. 3.
    We design an Anomalous cue Guided FAS feature extraction Network (AG-Net). The anomalous cue is introduced in different layers of the AG-Net via cross-attention modules to help obtain a more generalized FAS feature.
2 Related Work
@@
2.1 Face Anti-spoofing Methods
@@
Traditional FAS methods focus on designing hand-craft descriptors that reflect different representations between real and spoof faces. The most commonly used hand-craft descriptors in the field of computer vision, such as LBP [10, 11], HoG [28], SURF [12], SIFT [13], etc, are all leveraged as the representations for detecting the face presentation attacks. Although the hand-craft descriptors always have good interpretability, they struggle when dealing with more complex attack scenarios. Besides, deep neural networks are adopted to fix the FAS problem for their extraordinary feature learning capability. Commonly used neural networks like convolutional neural networks (CNNs)  [14, 29] and long short term memory (LSTM)  [15] are all introduced to the field of FAS and achieve great progress. With the help of the deep neural networks, other modalities like depth maps [16, 29] and reflection maps [30, 31] can be used jointly as auxiliary information for face liveness detection. However, though the above FAS methods achieve excellent performance on intra-dataset evaluations, they cannot well handle the data in unseen domains due to the limitation of the training FAS dataset diversity. Some researchers attempt to employ domain adaptation methods to improve the model’s performance in unseen domains. [32, 33, 34]. However, it remains uncertain which specific scenarios the model will encounter in real-world applications, and collecting data from all potential target datasets is challenging.
@@
To improve the performance of the FAS model in unseen domains, researchers have adopted domain generalization (DG) techniques that use multiple FAS datasets simultaneously to seek a discriminative domain-invariant feature invariant for liveness detection. Adversarial learning is mostly used in DG-based works to extract a common FAS feature space for all the domains  [17, 18]. A later work points out that the feature diversity differentiates between the real and spoof faces and adopts an asymmetric triplet loss function [19]. Another typical work concatenates the domain-invariant and the domain-related features from different images and uses contrastive learning to disentangle these two features [23]. Recent work finds it difficult to directly construct a domain-invariant feature space and proposes to align the live-to-spoof transition [35]. All existing FAS works based on domain generalization techniques achieve great improvement on several cross-dataset evaluation protocols. However, due to the lack of data diversity in existing FAS datasets, the capabilities of the previous FAS models are still constrained by the limited presentation attack types in the datasets.
2.2 Anomaly Detection Methods
@@
Another mainstream anomaly detection method is based on autoencoders. The autoencoder trained only on the normal samples can reconstruct the “normal” version image of the given input image. The difference between the generated “normal” image and the input image shows how close the input is to the normal distribution. A typical work uses a memory bank strategy to prevent the autoencoder from constructing an identical image as the input [41]. Subsequently, Generative adversarial networks (GANs) are employed to perform the function of the autoencoder, which generates samples that are more closely to the normal images [42, 43]. More recent anomaly detection works begin to explore the use of diffusion models as the autoencoder [44]. They add noise to images to blur the presence of the anomalous region and then use the diffusion model to remove the noise and generate high-quality normal images.
Compared to the autoencoders in the existing anomaly detection works that take the whole image as the input, the De-spoofing Face Generator (DFG) in our proposed method can generate a “real” face based solely on the input’s identity feature. On the one hand, restricting the input to only include identity feature helps prevent the DFG from generating identical images to the inputs, thereby increasing the distinction between an input spoofing face and its corresponding generated “real” face. On the other hand, the DFG can generate high-quality images with the help of large-scale real faces, reducing reconstruction errors when the input is a real face.
2.3 Diffusion Model
Diffusion model is a kind of multi-step generative model [45, 46], which gains widespread usage due to its ability to produce high-quality images. The early diffusion model is capable of generating from Gaussian noise unconditionally to obtain images that follow the distribution of the training dataset [45]. During the training process, the clean images $x_{0}$ are first progressively added noise for $n$ steps towards Gaussian noise $x_{n}$. Subsequently, given a step number $t$ and the corresponding noisy image $x_{t}$ as input, the model is trained to reconstruct the noisy image $x_{t-1}$ at step $t-1$. In the practical implementation of the paper, the target model output is the noise $\epsilon$ added to the image. The training objective can be simplified to:
$L_{DM}=\mathbb{E}_{x,\epsilon\in\mathcal{N}(0,1),t}[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}].$ (1)
@@ 2.3 Diffusion Model
$L_{CF}=\mathbb{E}_{x,c,\epsilon\in\mathcal{N}(0,1),t}[\|\epsilon-\epsilon_{\theta}(x_{t},t,c)\|_{2}^{2}].$ (2)
@@
$L_{LDM}=\mathbb{E}_{\mathcal{E}(x),y,\epsilon\in\mathcal{N}(0,1),t}[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau(y))\|_{2}^{2}],$ (3)
@@
where $\mathcal{E}$ is the encoder of a VQVAE which embeds the input $x$ into the latent space feature $z_{0}$ and then obtains the noisy $z_{t}$ via the diffusion process. $\tau$ is an encoder which projects the conditional input $y$ to an intermediate representation.
Considering the capability of diffusion models trained in the latent space to generate high-quality images, we choose it as the backbone of our De-spoofing Face Generator.
3 Method
3.1 Overview
Our method consists of two key modules, i.e., De-spoofing Face Generator (DFG) and Anomalous cue Guided FAS feature extraction Network (AG-Net), as illustrated in Figure 1 and Figure 2, respectively. As can be seen, we first train a De-spoofing Face Generator (DFG) on real faces. During training, a pre-trained Arcface [49] model is adopted to extract the face identity feature of the input, which is then taken as the conditional guidance of the DFG for face generation. Consequently, the trained DFG is capable of generating a “real” version face of the input when taking the input face identity feature as guidance. We then obtain an anomalous cue of the input image by calculating the residual between the generated “real” face image and the input. As the DFG is trained on real faces with rich data diversity, the anomalous cue can provide robust hints about the anomaly of the input under various scenarios.
For the training of a FAS model, we employ the DFG to generate the anomalous cue for each image in the FAS dataset as shown in Figure 2. The anomalous cue is then used to guide the training of a FAS feature extractor named Anomalous cue Guided FAS feature extraction Network (AG-Net). After being embedded by the anomalous cue encoder, the anomalous cue is introduced to the process of the FAS feature extraction at different layers via cross-attention modules. The learned FAS feature can benefit from the anomalous cue and generalize to unseen scenarios. Moreover, the AG-Net can be further combined with most existing domain generalization (DG) methods such as SSDG [19] to further enhance the generalization ability of the learned feature.
@@ 3.1 Overview
Figure 1: Structure of the proposed De-spoofing Face Generator (DFG). The DFG trained only on real faces can generate a corresponding “real” face of any given input face when taking the input’s identity feature as guidance.
3.2 De-spoofing Face Generator
Firstly, we employ a dataset that contains only real faces to train a De-spoofing Face Generator (DFG), as shown in Figure 1, aiming to obtain a model that can generate corresponding “real” faces when given any faces as input. We choose Latent Diffusion Model [48] as the backbone of DFG for its excellent generation performance.
During the training of Latent Diffusion Model [48], the text is encoded by a CLIP [50] model and used as the conditional guidance of the U-Net, which enables the model to generate images based on the text content. In the proposed DFG, we aim to generate a “real” version face of the same person in the input, so we take the face identity to be the conditional guidance of the generation process. Specifically, we use the face identity feature of the input face extracted by a face recognition model, Arcface [49], as the condition of the DFG. The last feature map before the global pooling layer of the Arcface feature extractor is chosen as the conditional input of the U-Net in DFG. The training objective for DFG is as follows:
$L_{DFG}=\mathbb{E}_{x,\mathcal{E}(x),\epsilon\in\mathcal{N}(0,1),t}[\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathcal{F}_{id}(x))\|_{2}^{2}],$ (4)
where $x$ is the input image, $\mathcal{E}$ is the VQVAE encoder which embeds the input $x$ into the latent code $z$ and then obtains the noisy $z_{t}$ via the diffusion process, $\epsilon$ is the noise added on the latent code, $\epsilon_{\theta}$ is the U-Net that is trained to predict the noise $\epsilon$, $t$ is the time step, and $\mathcal{F}_{id}$ is the Arcface feature extractor. The training objective of our DFG is similar to that of Latent Diffusion Models [48] in Eqn. 3, except that the conditional encoder is replaced with the Arcface feature extractor. During training, the parameters of the VQVAE encoder $\mathcal{E}$ and the Arcface feature extractor $\mathcal{F}_{id}$ are fixed, and only the parameters of the U-Net $\epsilon_{\theta}$ are optimized.
Once the U-Net is trained, the DFG can be used to generate “real” faces corresponding to the given faces. Specifically, the DFG first samples from a standard normal distribution to obtain the random noise $z_{T}$. Then, the U-Net takes the identity feature of the input face as the conditional input and progressively generates the latent code $z$ from the random noise $z_{T}$ in the latent space. Finally, the DFG uses the VQVAE decoder to obtain the “real” face from the latent code $z$ corresponding to the input.

### v2

# Generalized Face Liveness Detection via De-fake Face Generator

@@ Abstract
Previous Face Anti-spoofing (FAS) methods face the challenge of generalizing to unseen domains, mainly because most existing FAS datasets are relatively small and lack data diversity. Thanks to the development of face recognition in the past decade, numerous real face images are available publicly, which are however neglected previously by the existing literature. In this paper, we propose an Anomalous cue Guided FAS (AG-FAS) method, which can effectively leverage large-scale additional real faces for improving model generalization via a De-fake Face Generator (DFG). Specifically, by training on a large-scale real face only dataset, the generator obtains the knowledge of what a real face should be like, and thus has the capability of generating a “real” version of any input face image. Consequently, the difference between the input face and the generated “real” face can be treated as cues of attention for the fake feature learning. With the above ideas, an Off-real Attention Network (OA-Net) is proposed which allocates its attention to the spoof region of the input according to the anomalous cue. Extensive experiments on a total of nine public datasets show our method achieves state-of-the-art results under cross-domain evaluations with unseen scenarios and unknown presentation attacks. Besides, we provide theoretical analysis demonstrating the effectiveness of the proposed anomalous cues.
Index Terms:
@@ I Introduction
In today’s world, face recognition systems are extensively utilized for identity authentication in various scenarios. However, these systems are susceptible to presentation attacks, such as using a video or a photograph of a person to deceive the system, thereby posing a significant security threat. In order to address the security threat, significant efforts have been made in the field of Face Anti-Spoofing (FAS), also known as Presentation Attack Detection (PAD) according to the ISO definition [1], to defend against potential presentation attacks on face recognition systems. Researchers have constructed a diverse set of Face Anti-spoofing (FAS) datasets [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] and investigated the possibility of conducting face presentation attack detection on these datasets. Previous works have achieved excellent performance on each of these datasets under intra-dataset evaluation [12, 13, 14, 15, 16, 17, 18]. However, most existing FAS datasets are relatively small and lack data diversity. No individual dataset can cover all types of presentation attacks and various environments, which leads to domain bias for the FAS training. As a result, models trained on a single FAS dataset usually exhibit poor generalization on unseen FAS datasets [9, 19].
@@
To address this challenge, many efforts are devoted to domain generalization (DG) techniques to enhance the model’s generalization to unseen domains. Generally, the model is trained on multiple FAS datasets simultaneously to seek a discriminative domain-invariant feature for liveness detection. Some works resort to aligning the FAS features from different domains to make them domain indistinguishable [19, 20, 21, 22]. To further align the FAS feature, other interesting works focus on disentangling the domain-invariant and the domain-related features [23, 24, 25]. Besides, meta-learning is also utilized to help the model obtain the domain-invariant feature via simulating domain shift scenarios in the training stage [26, 27, 28, 29]. Although using domain generalization techniques relieves the lack of data diversity by training on multiple datasets, the capabilities of these methods are still constrained by the limited presentation attack types in the training datasets. Actually, we find numerous real faces can be easily achieved under various conditions and assume that effectively using these real faces is a key to improving the generalization of the FAS model. However, we find it non-trivial to utilize these data, as the straightforward addition of the large-scale real faces to the training data can lead to a performance decline in existing DG-based methods. We believe this might be attributed to the gap in the distribution between the additional real faces and the test FAS dataset, which introduces domain bias to the detection model.
@@
In this work, we propose an Anomalous cue Guided Face Anti-spoofing (AG-FAS) method that can effectively leverage additional real faces for improving cross-domain FAS. Specifically, we train a De-fake Face Generator (DFG) only on a large-scale real face dataset to gain the knowledge of what a real face should be like. The latent diffusion model based DFG can then remove the spoof trace of the input by adding noise during the forward process, i.e., the “de-fake” process, and restore a “real” version of the input via the generation process. As we utilize the face identity of the input to guide the generation process, we can adopt larger steps during the forward process of DFG to thoroughly remove the fake patterns, leading to a more “real” output. Consequently, we calculate the residual between the input and its “real” version, which is taken as the anomalous cue for the succeeding liveness detection. Then, we design an Off-real Attention Network (OA-Net) that utilizes the anomalous cue to allocate its attention toward the face region that deviates from the ”real”, i.e., the “off-real” region, achieving a more generalized FAS feature. Moreover, as a Plug-and-Play feature extraction module, the proposed OA-Net can be seamlessly tailored into most existing DG-based FAS methods to further improve the model performance across different domains. The experimental results show that the incorporation of this anomalous cue effectively enhances the performance of the model in unseen scenarios and unknown presentation attacks.
@@
  • We propose an Anomalous cue Guided Face Anti-spoofing (AG-FAS) method, which can effectively leverage a large scale of additional real faces for improving model generalization on cross-domain FAS tasks.
  • We train a De-fake Face Generator (DFG) that obtains an effective anomalous cue for liveness detection. With the help of the anomalous cue, the proposed Off-real Attention Network (OA-Net) can focus on the off-real region of the input, leading to a more generalized FAS feature.
  • We provide theoretical analysis on the distinguishability of anomalous cues between real and fake inputs, which further guarantees the feasibility of the proposed framework.
II Related Work
@@
II-A Face Anti-spoofing Methods
@@
Traditional FAS methods focus on designing hand-craft descriptors that reflect different representations between real and spoof faces. The most commonly used hand-craft descriptors in the field of computer vision, such as LBP [12, 13], HoG [30], SURF [14], SIFT [15], etc, are all leveraged as the representations for detecting the face presentation attacks. Although the hand-craft descriptors always have good interpretability, they struggle when dealing with more complex attack scenarios. Later, deep neural networks are adopted to fix the FAS problem for their extraordinary feature learning capability. Commonly used neural networks like convolutional neural networks (CNNs)  [16, 31] and long short term memory (LSTM)  [17] are all introduced to the field of FAS and achieve great progress. With the help of the deep neural networks, other modalities like depth maps [18, 31] and reflection maps [32, 33] can be used jointly as auxiliary information for face liveness detection. However, though the above FAS methods achieve excellent performance on intra-dataset evaluations, they cannot well handle the data in unseen domains due to the limitation of the training FAS dataset diversity. Some researchers attempt to employ domain adaptation methods to improve the model’s performance in unseen domains. [10, 34, 29]. However, it remains uncertain which specific scenarios the model will encounter in real-world applications, and collecting data from all potential target datasets is challenging.
@@
To improve the performance of the FAS model in unseen domains, researchers have adopted domain generalization (DG) techniques that use multiple FAS datasets simultaneously to seek a discriminative domain-invariant feature for liveness detection. Adversarial learning is mostly used in DG-based works to extract a common FAS feature space for all the domains  [19, 20]. A later work points out that the feature diversity differentiates between the real and spoof faces and adopts an asymmetric triplet loss function [21]. Another typical work concatenates the domain-invariant and the domain-related features from different images and uses contrastive learning to disentangle these two features [25]. Recent work finds it difficult to directly construct a domain-invariant feature space and proposes to align the live-to-spoof transition [35]. Many other works further enhance cross-domain generalization in FAS tasks by focusing on various aspects such as gradient regulation [36], hierarchical relations in samples [37], and the utilization of testing data [38]. With the rise of large vision-language models such as CLIP [39], a new line of works emerges that leverage text information to guide models in performing face anti-spoofing [40, 41]. All existing FAS works based on domain generalization techniques achieve great improvement on several cross-dataset evaluation protocols. However, due to the lack of data diversity in existing FAS datasets, the capabilities of the previous FAS models are still constrained by the limited presentation attack types in the datasets.
Some efforts are devoted to disentangling the attacks from the face by modeling the spoof traces [42, 43], leading to robust liveness detection. Nevertheless, effectively modeling all types of spoof traces is a non-trivial task. In contrast, our method focuses on modeling the distribution of real faces using achievable large-scale data, thereby allowing for better generalization to unseen scenarios and unknown presentation attacks.
II-B Anomaly Detection Methods
@@
Another mainstream anomaly detection method is based on image reconstruction. The model trained only on the normal samples is required to reconstruct the “normal” version image of the given input image. The difference between the generated “normal” image and the input image shows how close the input is to the normal distribution. A typical work based on autoencoder uses a memory bank strategy to prevent the model from constructing an identical image as the input [49]. Subsequently, generative adversarial networks (GANs) are employed to perform the function of image reconstruction [50, 51]. Since GAN training involves both an image encoder and decoder, GANs can similarly function as autoencoders to reconstruct the “normal” version of a given input image. Additionally, the adversarial learning in GANs allows for better reconstruction quality compared to most autoencoders, thereby improving anomaly detection performance. More recent anomaly detection works begin to explore the use of diffusion models for image reconstruction [52]. Unlike diffusion models used for unconditional image generation, here the model does not generate images directly from random noise. Instead, it first adds noise to the test image for a certain number of steps to remove the anomaly information. The model can then reconstruct a high-quality “normal” image based on the partially noised test image. This process, which resembles purifying the image, has also been used to defend against adversarial attacks in images [53].
The above idea of reconstructing normal samples and obtaining anomalous cues is also applied to the field of liveness detection. A pioneering work [54] attempts to utilize a de-spoofing diffusion model (DDM) to restore real faces from the input. The differences between the restored image and the input are used for detecting the presentation attacks. However, this work still faces many limitations, such as the trade-off between removing the spoof information and preserving the original face, as well as the effective utilization of the anomalous cues. Compared to DDM, we propose to utilize the face identity of the input for getting more distinguishable anomalous cues and introduce an attention-based feature extractor to better leverage the hints from the off-real region, resulting in a more generalized FAS feature.
II-C Diffusion Model
Diffusion model is a kind of multi-step generative model [55, 56], which gains widespread usage due to its ability to produce high-quality images. The early diffusion model is capable of generating from Gaussian noise unconditionally to obtain images that follow the distribution of the training dataset [55]. During the training process, the clean images $x_{0}$ are first progressively added noise for $n$ steps towards Gaussian noise $x_{n}$. Subsequently, given a step number $t$ and the corresponding noisy image $x_{t}$ as input, the model is trained to reconstruct the noisy image $x_{t-1}$ at step $t-1$. In the practical implementation of the paper [55], the target model output is the noise $\epsilon$ added to the image. The training objective can be simplified to:
$L_{DM}=\mathbb{E}_{x,\epsilon\in\mathcal{N}(0,1),t}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}.$ (1)
@@ II-C Diffusion Model
$L_{CDM}=\mathbb{E}_{x,c,\epsilon\in\mathcal{N}(0,1),t}\|\epsilon-\epsilon_{\theta}(x_{t},t,c)\|_{2}^{2}.$ (2)
@@
$L_{LDM}=\mathbb{E}_{E(x),y,\epsilon\in\mathcal{N}(0,1),t}\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau(y))\|_{2}^{2},$ (3)
@@
where $E$ is an encoder that embeds the input $x$ into the latent space feature $z_{0}$ and then obtains the noisy $z_{t}$ via the forward process. $\tau$ is another encoder which projects the conditional input $y$ to an intermediate representation.
@@
Figure 1: Structure of the proposed De-fake Face Generator (DFG). The DFG trained on a large-scale real face dataset can generate a corresponding “real” face of any given input face when taking the input’s identity feature $E_{id}(x)$ as guidance.
III Method
III-A Overview
Our method consists of two key modules, i.e., De-fake Face Generator (DFG) and Off-real Attention Network (OA-Net), as illustrated in Figure 1 and Figure 2, respectively. We first train a De-fake Face Generator (DFG) on a large-scale real face dataset. The trained DFG is capable of generating a “real” version face of the input when taking the input face identity feature $E_{id}(x)$ as guidance. We obtain an anomalous cue of the input image by calculating the residual between the generated “real” face image and the input. The anomalous cue is then used as a hint to guide the Off-real Attention Network (OA-Net) to explore more robust FAS features. As a Plug-and-Play feature extraction module, the OA-Net can be applied in most existing domain generalization (DG) methods such as SSDG [21] to further enhance the generalization ability of the learned feature.